Method of Elimination Word Problem
8. Mehrab works in a department store selling sports equipment. Baseball gloves cost $29 each and bats cost $14 each. One shift, he sells 28 items. His receipts total $647.
a) How many bats did Mehrab sell?
b) How many gloves did he sell?
Please i dont know how to do it. thank you very much!!!

Answer by tvandenberg(45)

G = Number of gloves
B = Number of bats
---
(a) B + G = 28
(b) 29*G + 14*B = 647
---
Solve (a) for B = 28 - G
Substitute into (b)
---
29*G + 14*(28-G) = 647
29*G + 392 - 14G = 647
15G = 255
G = 17
B = 28 - 17 = 11

Hope that helps!